The problem describes a debt to be amortized. (Round your answers to the nearest cent.) A man buys a boat for $310,000. He makes a $150,000 down payment and amortizes the rest of the purchase price with semiannual payments over the next 15 years. The interest rate on the debt is 13%, compounded semiannually. (a) Find the size of each payment. (b) Find the total amount paid for the purchase. (c) Find the total interest paid over the life of the loan.
Answer #1

Purchase Price of the boat = $310000

Down Payment = $150000

Balance Amount = $310000 - $150000 = $160000

Therefore,

P = $160000

r = 13% = 0.13

t = 15 years

n = 2  

a) Semiannual payment = rP/ n[1 - (1 + r/n)-nt]

= (0.13 * 160000)/ 2 [ 1 - (1 + 0.13/2)-2*15]

= (0.13 * 160000)/ 2 [ 1 - (1 + 0.13/2)-30]

= $12252.39

Size of each payment = $12252.39

b) Amount to be paid in 15 years = Semiannual Payment * n * t

= $12252.39 * 2 * 15

= $367571.70

Total amount paid for the purchase = Amount to be paid in 15 years + Down Payment

= $367571.70 + $150000

= $517571.70

c) Total interest paid = Amount to be paid in 15 years - P

= $367571.70 - $160000

= $207571.70

> B isn't working for me. This is the second time this site has old me to solve it this way and both times it's been wrong.

AlexConner0945 Mon, Nov 15, 2021 9:23 PM